ROSE'S GARDEN IS IN THE SHAPE OF A TRAPEZOID. IF THE HEIGHT OF THE TRAPEZOID IN 16M, ONE BASE IS 20M AND THE AREA IS 224M, FIND THE LENGTH OF THE OTHER BASE
Answer by tutorcecilia(2152)

A=1/2height(base#1 + base #2) [Use the formula for the area of a trapezoid]
224=(1/2)(16)(20+b) [Plug-in the values]
224=8(20+b)
224/8=8(20+b)/8
28=20+b [Solve for b (base #2)]
28-20=b
8=b
.
So, base #2 = 8M
.
Check by plugging all of the values back into the original equation.